OpenStudy (anonymous):

Write an equation of the line that is parallel to the given line y = x + 3 and the given point (5,0).

hartnn (hartnn):

slope of line y=x+3 can be found by comparing it with y=mx+c so, you get m=.... ?

OpenStudy (anonymous):

y = mx + c m = slope y = x + 3 y = 1x + 3 m = 1 Since the lines are parallel... m1 = m (5,0). Nw use point sSlope Form y - y1 = m ( x-x1)

OpenStudy (anonymous):

I got Y=5/0+3 but wasn't right

OpenStudy (anonymous):

y - 0 = 1 ( x-5) y = x - 5

OpenStudy (anonymous):

so I put y = x - 5?

hartnn (hartnn):

yes.